I haven't taken algebra in quite a few years. I'm trying to program a program to do a simple calculation. In order to do this, I have to know how to setup a word problem. My problem is to calculate how many weeks it would take to become a millionaire if I put 1 cent in my piggy bank one week, 2 cents the next week for a total of 3 cents, 3 cents the next week for a total of 6 cents, etc until I reach 1 million. I don't know how to set this up.
Answer by stanbon(75887)

My problem is to calculate how many weeks it would take to become a millionaire if I put 1 cent in my piggy bank one week, 2 cents the next week for a total of 3 cents, 3 cents the next week for a total of 6 cents, etc until I reach 1 million. I don't know how to set this up.
---
Series: 1+2+3+....+n = 100,000,000 cents
-----
Sum = (n/2)(1+n)
----
100,000,000 = (n/2)(1+n)
200,000,000 = n +n^2
-------
n^2+n-200,000,000 = 0
---
n = 14142 when rounded up
-------
That is the number of weeks you will need.
---
You will overshoot your goal by $25.77
but you will probably not be around to know it
unless you start very young.
============================================
Cheers,
Stan H.